# Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model

### v3

@@ 3.3 Inference-time Predictions
Figure 4 illustrates the inference process of KPDD. After fine-tuning, the process for solving a given question involves three main steps:
  1. 1.
    Core Question Extraction: First, we use the KPDD-CoT/PoT-core model to extract the core question from the original problem. This step isolates the essential part of the problem that needs to be addressed.
  2. 2.
    Problem-Solving Information Extraction: Next, the KPDD-CoT/PoT-info model extracts the relevant problem-solving information. This model identifies and lists the necessary context and data required to solve the core question.
  3. 3.
    Solution Generation: Finally, based on the original question, the core question, and the problem-solving information, the KPDD-CoT/PoT-solve model generates rationales in either CoT or PoT format to solve the original question. For KPDD-PoT, this involves generating Python code that can compute the answer.
This structured approach ensures that each model focuses on a specific aspect of the problem-solving process, leading to more accurate and reliable solutions.
@@ 4.1 Dataset
In our paper, we generate KPDD distillation datasets based on the GSM8K training set, which comprises diverse grade school math word problems (Cobbe et al., 2021). Then, we evaluate the mathematical reasoning performance of SLM on the GSM8K test set. Furthermore, to assess the transferability of SLMs’ mathematical reasoning capabilities, we evaluate SLM on several additional mathematical datasets. These datasets include ASDiv, which contains diverse math word problems (Miao et al., 2020), SVAMP, which features math word problems with varying structures (Patel et al., 2021), and MultiArith, which consists of arithmetic word problems (Roy and Roth, 2015). The statistics of these datasets are summarized in Table 1. This comprehensive evaluation approach ensures that the SLMs’ mathematical reasoning capabilities are thoroughly tested across a variety of problem types and structures, providing a robust assessment of their performance.
@@ 4.2 Implementation
In our paper, we use DeepSeek-V2 DeepSeek-AI (2024) as our teacher LLM to generate reasoning processes for questions. Specifically, we manually construct a demonstration set containing eight demonstrations. We then use this demonstration set to prompt DeepSeek-V2 to generate four reasoning paths for each question. Each reasoning path is subsequently filtered, resulting in the creation of a KPDD dataset. Next, we use FlanT5 models—Small (60M), Base (250M), and Large (760M) (Chung et al., 2022)—as our student LMs. By fine-tuning the FlanT5 models with the KPDD dataset, we aim to enhance their mathematical reasoning abilities. During the fine-tuning process, we set the learning rate to 5e-4, the batch size to 32, and the total number of training epochs to 10.
@@
Claude-2 (Anthropic, 2023) - 85.2 - - - 85.2
PaLM-2 (Anil et al., 2023) 540B 80.7 - - - 80.7
DeepSeek-V2 DeepSeek-AI (2024) 236B 92.2 - - - 92.2
@@
Llama-2 (Touvron et al., 2023) 7B 13.3 50.7 38.0 - 34
CodeLLaMA (Rozière et al., 2023) 7B 34.0 61.4 59.0 - 51.46
Platypus-2 (Lee et al., 2023) 7B 14.4 47.9 36.7 - 33
WizardMath (Luo et al., 2023) 7B 54.9 59.1 57.3 - 57.1
TORA (Gou et al., 2023) 7B 68.8 73.9 68.2 - 70.3
@@
(+) KPDD-CoT 21.75 22.51 19.1 35.5 24.71
(+) KPDD-PoT 46.32 59.92 61.6 87.5 63.83
Table 2: Overall test set performance.
@@ 4.3 Main Results
    KPDD-CoT Enhances Mathematical Reasoning: When FlanT5-small is used as the SLM, KPDD-CoT achieves an average accuracy improvement of 5.01% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-CoT yields an average accuracy improvement of 11.71%. For FlanT5-large, KPDD-CoT results in an average accuracy improvement of 15.51%. The experimental result demonstrates that KPDD-CoT can significantly enhance the mathematical reasoning performance of SLMs. We attribute this experimental result to the reason that baselines often encounters semantic misunderstanding errors that hinder the improvement of SLMs’ mathematical reasoning abilities. In contrast, KPDD-CoT employs extra SLMs to extract key points (including the core question and problem-solving information) of the question and uses these key points to guide the SLMs’ reasoning. This approach significantly reduces the semantic misunderstanding errors of CoTD, making KPDD-CoT better suited for improving the mathematical reasoning ability of SLMs.
@@
    KPDD-PoT Outperforms State-of-the-Art: When FlanT5-small is used as the SLM, KPDD-PoT achieves an average accuracy improvement of 32.18% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-PoT yields an average accuracy improvement of 48.25%. For FlanT5-large, KPDD-PoT results in an average accuracy improvement of 54.63%. The experimental result shows that KPDD-PoT make SLMs achieve state-of-the-art mathematical reasoning accuracy. Furthermore, KPDD-PoT’s accuracy is higher than that of KPDD-CoT, highlighting the advantage of rationales in PoT format in enhancing SLMs’ reasoning capabilities. Our analysis finds that the mathematical reasoning performance of CoTD is limited not only by semantic misunderstanding errors but also by calculation errors. PoTD converts rationales from CoT format into PoT format, formulating the reasoning process into a Python program and sending it to an extra Python interpreter to generate the final answer. This method transfers numerical computation from SLMs to a Python interpreter, avoiding calculation errors. Additionally, by extracting key points of the question, KPDD-PoT implicitly enhances the SLMs’ understanding of the question, thereby improving their overall mathematical reasoning capabilities.
@@
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
1 $\times$ $\times$ $\times$ 3.0 4.2 3.8 7.0 4.5
2 $\times$ $\times$ $\checkmark$ 8.71 9.2 8.2 10.33 9.11
3 $\checkmark$ $\times$ $\checkmark$ 9.02 9.25 8.9 11.5 9.66
4 $\times$ $\checkmark$ $\checkmark$ 8.87 9.73 8.9 11.0 9.59
5 $\checkmark$ $\checkmark$ $\checkmark$ 9.17 9.92 9.03 11.83 9.98
Table 3: Effect of Different Components in KPDD-CoT. We consider five different categories to analyse the effect of different components in KPDD-CoT. The experiment result shows that key points in questions can deepen SLMs’ understanding of the questions, and combining several key points can provide richer information, leading to further improvements in SLMs’ reasoning abilities.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
1 $\times$ $\times$ $\times$ 3.0 4.2 3.8 7.0 4.5
2 $\times$ $\times$ $\checkmark$ 19.40 44.32 40.6 45.33 37.41
3 $\checkmark$ $\times$ $\checkmark$ 23.19 45.89 44.1 53.33 41.62
4 $\times$ $\checkmark$ $\checkmark$ 25.39 46.85 44.6 57.33 43.54
5 $\checkmark$ $\checkmark$ $\checkmark$ 27.06 49.33 46.1 58.33 45.20
Table 4: Effect of Different Components in KPDD-PoT. We consider five different categories to analyse the effect of different components in KPDD-PoT. The experiment result shows that key points in questions can deepen SLMs’ understanding of the questions, and combining several key points can provide richer information, leading to further improvements in SLMs’ reasoning abilities.
@@ 4.4 Effect of Different Components in KPDD
In this subsection, we delve into the impact of various components within KPDD. We have considered five distinct categories, which include: 1. Original SLMs without any fine-tuning; 2. SLMs with original CoT/PoT distillation; 3. SLMs with core distillation combined with CoT/PoT distillation; 4. SLMs with problem-solving information distillation combined with CoT/PoT distillation; 5. SLMs with KPDD. For each of the latter four categories, we have constructed corresponding reasoning datasets, each containing a single reasoning path per question. Following this, we have utilized FlanT5-base as our foundation for SLMs, and we have fine-tuned these models using the aforementioned reasoning datasets. To evaluate the reasoning capabilities of these SLMs, we have tested them on the GSM8K test dataset, as well as on the ASDiv, SVAMP, and MultiArith datasets.
@@
Tables 3 and 4 present the results of our experiments, from which we make several observations: (1) We observe a significant performance improvement in Category 2 compared to original SLMs. Specifically, under CoT reasoning, Category 2 achieves an average accuracy gain of 4.61% across multiple datasets, while under PoT reasoning, it achieves a substantial average accuracy improvement of 32.91%. These experimental results indicate that CoTD and PoTD can markedly enhance the mathematical reasoning ability of SLMs. (2) We find that Categories 3 and 4 exhibit a further performance increase relative to Category 2. Specifically, in the context of CoT reasoning, Categories 3 and 4 achieve average accuracy gains of 0.55% and 0.45% respectively over Category 2 across multiple datasets. Under PoT reasoning, the gains are more pronounced with Categories 3 and 4 achieving average accuracy improvements of 4.21% and 6.13% respectively. This suggests that SLMs can deepen their understanding of questions by focusing on key points, thereby further enhancing their mathematical reasoning ability. (3) In Category 5, we combine the core questions with the problem-solving information to guide SLMs in addressing the questions. The results are promising: Category 5 achieves an average accuracy of 9.98% under CoT reasoning and a remarkable 45.20% under PoT reasoning across multiple datasets. This indicates that key points in questions play a crucial role in boosting the reasoning capabilities of SLMs, and that combining several key points provides richer information, leading to further improvements in their reasoning abilities.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
$\mathrm{I}$ 1* 1 1 7.88 4.72 5.4 10.66 7.16
$\mathrm{II}$ 1 1 2 9.09 9.44 8.2 11.33 9.51
$\mathrm{III}$ 1 2 2 8.41 7.72 6.7 11.24 8.51
$\mathrm{IV}$ 2 1 2 7.80 7.58 7.1 11.16 8.41
$\mathrm{V}$ 1 2 3 9.17 9.92 9.03 11.83 9.98
    The index of SLM.
Table 5: Effect of SLM Quantity in KPDD-CoT. We consider five different categories to analyse the effect of SLM quantity in KPDD-CoT. The experimental results show that for KPDD-CoT, using a separate SLM for each component is necessary to maximize the reasoning performance of KPDD-CoT.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
$\mathrm{I}$ 1* 1 1 24.18 44.32 41.19 48.66 39.61
$\mathrm{II}$ 1 1 2 26.0 42.69 42.69 55.83 41.80
$\mathrm{III}$ 1 2 2 24.79 46.37 40.6 49.16 40.23
$\mathrm{IV}$ 2 1 2 24.63 45.37 41.3 49.33 40.15
$\mathrm{V}$ 1 2 3 27.06 49.33 46.1 58.33 45.20
    The index of SLM.
Table 6: Effect of SLM Quantity in KPDD-PoT. We consider five different categories to analyse the effect of SLM quantity in KPDD-PoT. The experimental results show that for KPDD-PoT, using a separate SLM for each component is necessary to maximize the reasoning performance of KPDD-PoT.
4.5 Effect of SLM Quantity in KPDD
In this subsection, we investigate the impact of SLM Quantity in KPDD. We consider five distinct categories: $\mathrm{I}$. Using one SLM to simultaneously extract the core question and problem-solving information, and solve the original question; $\mathrm{II}$. Using one SLM to extract the core question and problem-solving information, and another SLM to solve the original question; $\mathrm{III}$. Using one SLM to extract the core question, another SLM to extract the problem-solving information, and a third SLM to solve the original question; $\mathrm{IV}$. Using one SLM to extract the problem-solving information, another SLM to extract the core question, and both to solve the original question; $\mathrm{V}$. Using one SLM to extract the core question, another SLM to extract the problem-solving information, and a third SLM to solve the original question. For each category, we create corresponding reasoning datasets, each containing a single reasoning path per question. We utilize FlanT5-base as our base SLMs, fine-tuning them on these reasoning datasets. To assess their reasoning capabilities, we evaluate these SLMs on the GSM8K test dataset, as well as on the ASDiv, SVAMP, and MultiArith datasets.
Table 5 and 6 present the results of our experiments, from which we make several observations: (1) Compared to other categories, Category $\mathrm{I}$ performed worse. For KPDD-CoT, Category $\mathrm{I}$ achieved an average accuracy of 7.16% across multiple datasets, while for KPDD-PoT, it achieved an average accuracy of 39.61%. This suggests that the limited model size of a single SLM hinders its performance across multiple tasks. (2) Category $\mathrm{II}$ outperformed Categories $\mathrm{III}$ and $\mathrm{IV}$ in reasoning performance. For KPDD-CoT, Category $\mathrm{II}$ achieved an average accuracy of 9.51% across multiple datasets, while for KPDD-PoT, it achieved an average accuracy of 41.80%. We attribute this result to the importance of the KPDD-CoT/PoT-solve component, where using a single SLM for this phase yields the best reasoning performance. (3) For KPDD-CoT, Category $\mathrm{V}$ achieved an average accuracy of 9.98% across multiple datasets, while for KPDD-PoT, it achieved an average accuracy of 45.20%. This is the highest reasoning performance among all categories, indicating that our approach of using a separate SLM for each component maximizes the performance of each component, thereby maximizing the reasoning performance of KPDD.
@@ 4.5 Effect of SLM Quantity in KPDD
Figure 5: Effect of Reasoning Paths. We fine-tune CodeT5-Base with different reasoning paths to analyse the effect of reasoning paths. The experiment results shows that diverse reasoning paths can improve SLMs’ reasoning performance.
@@ 4.6 Diverse Reasoning Paths Improve SLMs’ Reasoning Performance
Figure 5 reveals the relevant experimental results. As the number of reasoning paths in the distillation dataset increases, the SLM tends to achieve better performance. For example, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing a single reasoning path, it achieved an accuracy of 9.17% on the GSM8K test dataset, 9.92% on ASDiv, 9.03% on SVAMP, and 11.83% on MultiArith. However, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing four reasoning paths, it achieved accuracies of 14.63% on the GSM8K test dataset, 14.93% on ASDiv, 13.8% on SVAMP, and 21.5% on MultiArith. Similarly, when fine-tuning SLMs with KPDD-PoT datasets containing different reasoning paths, the experimental results were consistent with those of KPDD-CoT. These results indicate that a mathematical reasoning dataset with diverse reasoning paths can effectively enhance the mathematical reasoning performance of SLMs.
@@ 4.7 Error Analysis
In this subsection, our aim is to verify whether KPDD can indeed reduce semantic misunderstanding errors. KPDD-PoT implicitly includes the reasoning process within its rationales, making it challenging to conduct error analysis on rationales in PoT format. Conversely, rationales in CoT format explicitly contain the reasoning steps, allowing us to clearly understand how the SLM solves the questions step by step, thus facilitating error analysis. Therefore, in this part, we focus on error analysis for rationales in CoT format. To achieve our goal, we randomly sample 100 examples from GSM8K/SVAMP and perform error analysis on the questions with incorrect answers. For a better understanding of KPDD’s effect, we also consider three other scenarios: (1) vanilla CoTD, (2) reasoning that combines vanilla CoTD and core question extraction, and (3) reasoning that combines vanilla CoTD and problem-solving information extraction. Furthermore, to simplify our analysis, we use flanT5-base as our SLMs, and the corresponding reasoning datasets still contain a single reasoning path per question.
The detailed quantitative results are illustrated in Figure 6. By analyzing the experimental results, we found that: (1) Combination of Multiple Errors in SLMs: SLMs tend to exhibit combinations of multiple errors, with calculation errors having the most significant impact on reasoning performance. Specifically, vanilla CoTD on the GSM8K dataset showed 51 understanding errors, 79 calculation errors, and 34 step missing errors, resulting in a total of 164 errors. This number far exceeds the original number of problems, with calculation errors outnumbering other types of errors. Similar results were observed in the SVAMP dataset. This explains why PoTD achieves better reasoning performance than CoTD: PoTD converts vanilla rationales into Python programs, delegating the calculation process to an external Python interpreter to avoid calculation errors. (2) Reduction of Understanding Errors with Key Points: Introducing key points of the original questions effectively reduces understanding errors. Specifically, when core questions were introduced in vanilla CoTD, the number of understanding errors on the GSM8K dataset decreased to 50, and on the SVAMP dataset, it decreased to 53. When problem-solving information was introduced in vanilla CoTD, the number of understanding errors decreased to 48 on GSM8K and to 51 on SVAMP. These results indicate that key points of the original questions help SLMs better understand the questions, thereby reducing understanding errors and improving reasoning performance. (3) Further Reduction of Understanding Errors with Multiple Key Points: Combining multiple key points can further reduce understanding errors. Specifically, KPDD reduced the number of understanding errors to 46 on GSM8K and to 50 on SVAMP. This suggests that KPDD’s method of integrating multiple key points can deepen SLMs’ understanding of the original questions, further reducing understanding errors and enhancing reasoning performance.

### v4

@@ 3.3 Inference-time Predictions
After fine-tuning, the process for solving a question involves two steps. Firstly, we use the KPDD-CoT/PoT-Key SLM to extract the core question and the relevant problem-solving information from the original question. This SLM identifies and outlines the key points necessary for KPDD-CoT/PoT-Solve, enhancing the understanding of the original question. Then, based on the original question, the key points, the KPDD-CoT/PoT-Solve SLM generates rationales in either CoT or PoT format to solve the original question. For KPDD-PoT, this requires an additional Python interpreter to execute the generated Python program and obtain the final answer. The approach ensures that each SLM focuses on a specific aspect of the problem-solving process, leading to more accurate and reliable solutions.
@@ 4.1 Dataset
In our paper, we generate KPDD distillation datasets based on the GSM8K training set, which comprises diverse grade school math word problems Cobbe et al. (2021). Then, we evaluate the mathematical reasoning performance of SLM on the GSM8K test set. Furthermore, to assess the transferability of SLMs’ mathematical reasoning capabilities, we evaluate SLM on several additional mathematical datasets. These datasets include ASDiv, which contains diverse math word problems Miao et al. (2020), SVAMP, which features math word problems with varying structures Patel et al. (2021), and MultiArith, which consists of arithmetic word problems Roy and Roth (2015). The statistics of these datasets are summarized in Table 1. This comprehensive evaluation approach ensures that the SLMs’ mathematical reasoning capabilities are thoroughly tested across a diverse range of both in-domain and out-of-domain questions, providing a robust assessment of their performance.
@@ 4.2 Implementation
In our paper, we use ChatGPT as our teacher LLM to generate reasoning processes for questions. Specifically, we manually construct a demonstration set containing eight demonstrations. We then use this demonstration set to prompt ChatGPT to generate four reasoning paths for each question. Each reasoning path is subsequently filtered, resulting in the creation of a KPDD dataset. Next, we use FlanT5 models—Small (60M), Base (250M), and Large (760M) Chung et al. (2022)—as our student LMs. By fine-tuning the FlanT5 models with the KPDD dataset, we aim to enhance their mathematical reasoning abilities. During the fine-tuning process, we set the learning rate to 5e-4, the batch size to 32, and the total number of training epochs to 10. The input and output lengths for KPDD-CoT/PoT-Key are set to 256 and 256, respectively, while the input and output lengths for KPDD-CoT/PoT-Solve are set to 512 and 256, respectively.
@@
Claude-2 Anthropic (2023) - 85.2 - - - 85.2
PaLM-2 Anil et al. (2023) 540B 80.7 - - - 80.7
@@
Llama-2 Touvron et al. (2023) 7B 13.3 50.7 38.0 - 34
CodeLLaMA Rozière et al. (2023) 7B 34.0 61.4 59.0 - 51.46
Platypus-2 Lee et al. (2023) 7B 14.4 47.9 36.7 - 33
WizardMath Luo et al. (2023) 7B 54.9 59.1 57.3 - 57.1
TORA Gou et al. (2023) 7B 68.8 73.9 68.2 - 70.3
@@
(+) KPDD-CoT 20.92 27.05 21.5 40 27.36
(+) KPDD-PoT 49.05 60.16 58.09 90.33 64.4
Table 2: Overall Test Sets Performance.
@@ 4.3 Main Results
    KPDD-CoT Enhances Mathematical Reasoning: When FlanT5-small is used as the SLM, KPDD-CoT achieves an average accuracy improvement of 7.35% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-CoT yields an average accuracy improvement of 13.34%. For FlanT5-large, KPDD-CoT results in an average accuracy improvement of 18.16%. The experimental result demonstrates that KPDD-CoT can significantly enhance the mathematical reasoning performance of SLMs. We attribute this experimental result to the reason that baselines often encounters semantic misunderstanding errors that hinder the improvement of SLMs’ mathematical reasoning abilities. In contrast, KPDD-CoT employs extra SLMs to extract key points (including the core question and problem-solving information) of the question and uses these key points to guide the SLMs’ reasoning. This approach significantly reduces the semantic misunderstanding errors of CoTD, making KPDD-CoT better suited for improving the mathematical reasoning ability of SLMs.
@@
    KPDD-PoT Outperforms State-of-the-Art: When using FlanT5-small as the SLM, KPDD-PoT improves average accuracy by 34.32% in mathematical reasoning tasks. With FlanT5-base, it increases by 49.39%, and with FlanT5-large, by 55.2%. This shows that KPDD-PoT helps SLMs achieve state-of-the-art reasoning accuracy. It also outperforms KPDD-CoT, demonstrating the benefits of PoT-format rationales in improving reasoning. This is because that CoTD’s performance is limited not only by semantic misunderstanding errors but also by calculation errors. PoTD turns CoT rationales into Python code and uses a Python interpreter for final answers, avoiding calculation errors. KPDD-PoT also enhances question understanding, improving overall reasoning performance.
@@
Reasoning Format Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
CoT 1 $\times$ $\times$ $\times$ 3.0 4.2 3.8 7.0 4.5
2 $\times$ $\times$ $\checkmark$ 7.05 10.54 7.3 11.5 9.09
3 $\checkmark$ $\times$ $\checkmark$ 7.35 12.4 8.2 11.16 9.77
4 $\times$ $\checkmark$ $\checkmark$ 7.73 12.78 7.7 11.83 10.01
5 $\checkmark$ $\checkmark$ $\checkmark$ 8.64 12.92 8.7 13.83 11.02
PoT 1 $\times$ $\times$ $\times$ 3.0 4.2 3.8 7.0 4.5
2 $\times$ $\times$ $\checkmark$ 20.09 42.84 36.7 44.83 36.11
3 $\checkmark$ $\times$ $\checkmark$ 25.17 46.42 39.3 51.33 40.55
4 $\times$ $\checkmark$ $\checkmark$ 25.32 46.94 40.2 51.66 41.03
5 $\checkmark$ $\checkmark$ $\checkmark$ 26.08 47.32 41.5 53.83 42.18
Table 3: Effect of Different Components in KPDD. We consider five different categories to analyse the effect of different components in KPDD. The experiment result shows that key points in questions can deepen SLMs’ understanding of the questions, and combining several key points can provide richer information, leading to further improvements in SLMs’ reasoning abilities.
@@ 4.4 Effect of Different Components in KPDD
In this subsection, we delve into the impact of various components within KPDD. We have considered five distinct categories, which include: 1. Original SLMs without any fine-tuning; 2. SLMs with original CoT/PoT distillation; 3. SLMs with core distillation combined with CoT/PoT distillation; 4. SLMs with problem-solving information distillation combined with CoT/PoT distillation; 5. SLMs with KPDD. For each of the latter four categories, we have constructed corresponding distillation datasets, each containing a single reasoning path per question. Following this, we have utilized FlanT5-base as our foundation for SLMs, and we have fine-tuned these models using the aforementioned distillation datasets. To evaluate the reasoning capabilities of these SLMs, we have tested them on the GSM8K test dataset, as well as on the ASDiv, SVAMP, and MultiArith datasets.
@@
Tables 3 present the results of our experiments, from which we make several observations: (1) We observe a significant performance improvement in Category 2 compared to original SLMs. Specifically, under CoT reasoning, Category 2 achieves an average accuracy gain of 4.59% across multiple datasets, while under PoT reasoning, it achieves a substantial average accuracy improvement of 31.61%. These experimental results indicate that CoTD and PoTD can markedly enhance the mathematical reasoning ability of SLMs. (2) We find that Categories 3 and 4 exhibit a further performance increase relative to Category 2. Specifically, in the context of CoT reasoning, Categories 3 and 4 achieve average accuracy gains of 0.77% and 0.92% respectively over Category 2 across multiple datasets. Under PoT reasoning, the gains are more pronounced with Categories 3 and 4 achieving average accuracy improvements of 4.44% and 4.92% respectively. This suggests that SLMs can deepen their understanding of questions by focusing on key points, thereby further enhancing their mathematical reasoning ability. (3) In Category 5, we combine the core questions with the problem-solving information to guide SLMs in addressing the questions. The results are promising: Category 5 achieves an average accuracy of 11.02% under CoT reasoning and a remarkable 42.18% under PoT reasoning across multiple datasets. This indicates that key points in questions play a crucial role in boosting the reasoning capabilities of SLMs, and that combining several key points provides richer information, leading to further improvements in their reasoning abilities.
@@
Figure 3: Effect of Reasoning Paths. We fine-tune FlanT5-Base with different reasoning paths to analyse the effect of reasoning paths. The experiment results shows that diverse reasoning paths can improve SLMs’ reasoning performance.
@@ 4.5 Diverse Reasoning Paths Improve SLMs’ Reasoning Performance
Figure 3 reveals the relevant experimental results. As the number of reasoning paths in the distillation dataset increases, the SLM tends to achieve better performance. For example, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing a single reasoning path, it achieved an accuracy of 8.64% on the GSM8K test dataset, 12.92% on ASDiv, 8.7% on SVAMP, and 13.83% on MultiArith. However, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing four reasoning paths, it achieved accuracies of 14.32% on the GSM8K test dataset, 19.32% on ASDiv, 13.9% on SVAMP, and 23.83% on MultiArith. Similarly, when fine-tuning SLMs with KPDD-PoT datasets containing different reasoning paths, the experimental results were consistent with those of KPDD-CoT. These results indicate that a mathematical reasoning dataset with diverse reasoning paths can effectively enhance the mathematical reasoning performance of SLMs.
@@ 4.6 Error Analysis
In this subsection, our aim is to verify whether KPDD can indeed reduce semantic misunderstanding errors. KPDD-PoT implicitly includes the reasoning process within its rationales, making it challenging to conduct error analysis on rationales in PoT format. Conversely, rationales in CoT format explicitly contain the reasoning steps, allowing us to clearly understand how the SLM solves the questions step by step, thus facilitating error analysis. Therefore, in this part, we focus on error analysis for rationales in CoT format. To achieve our goal, we randomly sample 100 examples from GSM8K/ASDiV and perform error analysis on the questions with incorrect answers. For a better understanding of KPDD’s effect, we also consider three other scenarios: (1) vanilla CoTD, (2) reasoning that combines CoTD and core question extraction, and (3) reasoning that combines CoTD and problem-solving information extraction. Furthermore, to simplify our analysis, we use FlanT5-base as our SLMs, and the corresponding reasoning datasets still contain a single reasoning path per question.
The detailed quantitative results are illustrated in Figure 4. By analyzing the experimental results, we found that: (1) Calculation Errors and Understanding Errors are The Major Errors: The experiment results indicate that calculation errors and misunderstanding errors are the primary factors limiting CoTD’s mathematical reasoning performance. Specifically, vanilla CoTD on the GSM8K dataset showed 36 understanding errors, 45 calculation errors, and 11 step missing errors. Similar results were observed in the ASDiV dataset. This explains why PoTD achieves better reasoning performance than CoTD: PoTD converts vanilla rationales into Python programs, delegating the calculation process to an external Python interpreter to avoid calculation errors. (2) Reduction of Understanding Errors with Key Points: Introducing key points of the original questions effectively reduces understanding errors. Specifically, when core questions were introduced in vanilla CoTD, the number of understanding errors on the GSM8K dataset decreased to 33, and on the ASDiV dataset, it decreased to 22. When problem-solving information was introduced in vanilla CoTD, the number of understanding errors decreased to 32 on GSM8K and to 21 on ASDiV. These results indicate that key points of the original questions help SLMs better understand the questions, thereby reducing understanding errors and improving reasoning performance. (3) Further Reduction of Understanding Errors with Multiple Key Points: Combining multiple key points can further reduce understanding errors. Specifically, KPDD reduced the number of understanding errors to 30 on GSM8K and to 16 on ASDiV. This suggests that KPDD’s method of integrating multiple key points can deepen SLMs’ understanding of the original questions, further reducing understanding errors and enhancing reasoning performance.